An airplane circle the earth at 0.6c.How many years must elaspe before a clock in the plane and the one on the ground differ by 1S
1
Expert's answer
2021-07-29T16:42:02-0400

Gives

v=0.6c

Now

We know that

"∆x=x_0\sqrt{1-\frac{v^2}{c^2}}"

"∆x=x_0\sqrt{1-\frac{(0.6c)^2}{c^2}}"

"∆x= x_0\sqrt{1-0.36}=0.8x_0"

Time

"∆t=\frac{∆x}{v}"


"∆t=\frac{0.8x_0}{0.6c}=\frac{0.8 x_0}{0.6\times3\times10^8\times60\times60\times365}=3.382\times10^{-15} years"